Smith Distributors, Inc., supplies ice cream shops with various toppings for making sundaes. On November 17, 2021, a fire resulted in the loss of all of the toppings stored in one section of the warehouse. The company must provide its insurance company with an estimate of the amount of inventory lost. The following information is available from the company's accounting records:

Fruit
Toppings
Marshmallow
Toppings
Chocolate
Toppings
Inventory, January 1, 2021 $ 15,000 $ 6,500 $ 2,500
Net purchases through Nov. 17 125,000 31,000 11,500
Net sales through Nov. 17 175,000 50,000 19,500
Historical gross profit ratio 30 % 35 % 35 %


Required:
1. Calculate the estimated cost of each of the toppings lost in the fire.

Estimated cost of
Toppings lost inventory
Fruit
Marshmallow
Chocolate
Answer #1

1) calculate estimated cost of lost in the fire

Fruit Marshmallow Chocolate
Beginning inventory 15000 6500 2500
Purchase 125000 31000 11500
Cost of goods available for sale 140000 37500 14000
Less: Cost of goods sold 175000*70% = -122500 50000*65% = -32500 19500*65% = -12675
Inventory lost in fire 17500 5000 1325